# Comparative Analysis of Weather-Based Indexes and the Actuaries Climate IndexTM for Crop Yield Prediction

### v1

@@ 5 Concluding Remarks
This study examines the Actuaries Climate IndexTM within the agricultural sector and compares its effectiveness with widely recognized weather-based indexes. The focus is on evaluating the predictive capabilities of both ACI components and traditional weather-based indexes in predicting three selected crop yields. By fitting 22 distinct models, each constructed with different climate indexes and their combinations, this paper also investigates the individual contribution of each weather index to model performance. Based on high goodness-of-fit test results and low MAPE values, climate indexes exhibit strong predictive power in modelling crop yields. The key insights drawn from the analyses can be summarized as follows:
@@
    The four principal components derived through FPCA are sufficiently effective in capturing the underlying patterns in crop yield variations.
@@
    ACI components and their WBI counterparts exhibit strong correlation, and deliver similar predictive performance in predicting crop yields, especially with machine learning algorithms.
@@
    The inclusion of indexes representing additional weather conditions—such as wind speed, and sea-level changes—alongside temperature and precipitation, significantly improves model performance.
@@
The outcomes suggest that the ACI framework has a potential to serve as a valuable climate risk indicator not only for the agricultural sector but also for the finance and insurance industries, where climate-related metrics are becoming increasingly critical for risk assessment and decision-making.
Despite the promising performance metrics achieved, the study can be further improved in several aspects. One notable limitation is that regional crop yields were assessed by averaging the annual yields of the states within each region. This approach overlooks the fact that some states may have a greater influence on regional agricultural output due to their size or productivity levels. Moreover, the selection of the best-performing models for each crop is based on regional average results, which may not accurately reflect optimal performance at the regional level. Consequently, identifying and evaluating the best models for each specific region or state would enhance the precision and applicability of the findings, particularly for targeted applications in the insurance sector.
@@ Appendix A Preliminaries
In this section, we provide an overview of generalized linear models (GLMs) and generalized additive models (GAMs), which belong to the class of generalized statistical models, as well as Extreme Gradient Boosting (XGB) and Light Gradient Boosting Machine (LGBM), which are classified as machine learning algorithms. Furthermore, we review principal component analysis (PCA) and functional principal component analysis (FPCA), two dimensionality reduction techniques used to preprocess and simplify data sets before being incorporated into the models.
@@ A.1 Generalized Statistical Models
where $\beta_{0}$ is the intercept term, and $f_{j}(X_{j})$ $(j=1,\dots,p)$ are smooth functions that model the non-linear effects of the predictors $X_{j}$ on the response variable. The flexibility of GAMs allow them to capture complex relationships while retaining interpretability, making it an effective tool for modeling real-world problems with non-linear dependencies.
@@ A.2 Machine Learning Algorithms
Here, $\ell$ denotes a differentiable loss function, $\hat{y}_{i}^{(t-1)}$ is the prediction at iteration $t-1$, $\Omega(f_{t})$ penalizes the complexity of each tree $f_{t}$, with $T$ being the number of leaves and $\mathbf{w}$ the vector of leaf weights.
@@ A.3 Dimensionality Reduction in Modeling
Given a data matrix $X$ of size $n\times p$ where $n$ represents the number of observations and $p$ is the number of variables, each element in $X$ can be standardized around its mean value, normalized by its standard deviation, thus obtaining a matrix of $Z$. PCA seeks to find the principal components by calculating the eigenvectors and eigenvalues of the covariance matrix, $\Sigma_{p\times p}$, given by,
@@
[…]
Appendix B Tables and Figures
@@
Group-1 CDD Mean ± SD 504 ± 96 243 ± 72 617 ± 109 1175 ± 109 794 ± 90 888 ± 100
Min. 326 84 356 928 526 697
Max. 703 427 839 1464 1016 1086
T90 Mean ± SD 0.40 ± 1.20 0.49 ± 1.31 0.01 ± 0.89 0.54 ± 1.21 0.54 ± 1.42 1.06 ± 1.69
Min. -1.48 -1.65 -1.32 -1.25 -1.61 -1.91
Max. 3.57 5.15 3.57 4.49 5.32 5.13
Group-2 HDD Mean ± SD 3274 ± 275 2727 ± 184 3658 ± 326 1877 ± 236 3346 ± 269 2375 ± 166
Min. 2715 2285 3002 1415 2797 2024
Max. 3832 3237 4382 2452 4208 2799
T10 Mean ± SD -0.46 ± 1.05 -0.15 ± 0.89 -0.30 ± 0.96 -0.44 ± 0.98 -0.28 ± 0.93 -0.31 ± 0.88
@@
Group-3 PRE Mean ± SD 11.29 ± 2.21 8.31 ± 1.86 10.07 ± 1.73 13.48 ± 2.61 7.02 ± 1.19 3.61 ± 1.04
Min. 6.79 3.88 6.10 7.49 4.49 1.54
Max. 19.04 13.49 13.82 10.21 10.21 7.02
P Mean ± SD 0.10 ± 1.00 0.58 ± 1.16 0.26 ± 1.03 -0.02 ± 0.94 0.33 ± 1.39 0.19 ± 0.97
Min. -2.05 -1.43 -2.13 -2.07 -1.65 -1.92
Max. 3.48 2.85 2.95 1.85 5.27 2.93
@@
(a) Group-1
(b) Group-2
(c) Group-3
Figure B1: Group indexes of each region over time
(a) Group-1
(b) Group-2
(c) Group-3
Figure B2: CDF plots of group indexes for each region
[…]
Crop Region Model
@@
Corn (Grain) CEA 0.174 0.276 0.345 0.197 0.493 0.466 0.458 0.122 0.251 0.340 0.347 0.397 0.417 0.374 0.421 0.344 0.632 0.431 0.608 0.608 0.604 0.428
CWP 0.297 0.292 0.182 0.319 0.240 0.391 0.361 0.313 0.282 0.502 0.322 0.552 0.565 0.593 0.538 0.493 0.452 0.666 0.549 0.504 0.303 0.328
MID 0.228 0.236 0.249 0.184 0.336 0.291 0.312 0.112 0.101 0.312 0.022 0.322 0.297 0.307 0.338 0.237 0.307 0.292 0.338 0.237 0.292 0.156
SEA 0.295 0.334 0.113 0.292 0.260 0.280 0.342 0.172 0.214 0.198 0.195 0.185 0.312 0.311 0.231 0.292 0.796 0.220 0.751 0.704 0.687 0.520
SPL 0.251 0.198 0.105 0.224 0.193 0.212 0.299 0.130 0.128 0.280 0.120 0.286 0.239 0.306 0.337 0.381 0.755 0.386 0.721 0.734 0.736 0.533
SWP 0.481 0.441 0.026 0.499 0.481 0.479 0.533 0.455 0.412 0.121 0.486 0.489 0.516 0.542 0.487 0.523 0.504 0.524 0.538 0.519 0.466 0.452
Wheat (Winter) CEA 0.328 0.473 0.065 0.486 0.308 0.440 0.414 0.274 0.444 0.122 0.338 0.334 0.368 0.378 0.484 0.406 0.689 0.478 0.621 0.668 0.674 0.655
CWP 0.125 0.257 0.216 0.160 0.226 0.390 0.248 0.231 0.270 0.567 0.281 0.661 0.646 0.626 0.581 0.483 0.491 0.487 0.517 0.486 0.290 0.321
MID 0.202 0.234 0.117 0.284 0.220 0.266 0.281 0.102 0.123 0.159 0.128 0.152 0.190 0.198 0.359 0.262 0.198 0.305 0.359 0.262 0.305 0.174
SEA 0.295 0.283 0.067 0.271 0.193 0.223 0.242 0.227 0.223 0.111 0.235 0.178 0.111 0.122 0.241 0.117 0.643 0.195 0.606 0.631 0.578 0.570
SPL 0.281 0.200 0.203 0.277 0.361 0.358 0.387 0.141 0.201 0.308 0.128 0.341 0.377 0.365 0.280 0.324 0.735 0.306 0.774 0.741 0.727 0.528
SWP 0.341 0.326 0.038 0.330 0.382 0.394 0.423 0.317 0.300 0.167 0.340 0.492 0.404 0.442 0.435 0.403 0.398 0.426 0.433 0.403 0.387 0.384
Soybeans CEA 0.201 0.328 0.368 0.256 0.542 0.565 0.553 0.167 0.285 0.332 0.375 0.434 0.440 0.430 0.387 0.438 0.614 0.437 0.597 0.624 0.596 0.462
MID 0.281 0.282 0.255 0.250 0.391 0.362 0.416 0.094 0.152 0.336 0.052 0.346 0.345 0.366 0.389 0.326 0.366 0.344 0.389 0.326 0.344 0.148
SEA 0.231 0.240 0.248 0.196 0.318 0.395 0.419 0.096 0.149 0.284 0.130 0.193 0.384 0.392 0.285 0.316 0.664 0.259 0.666 0.600 0.572 0.390
SPL 0.285 0.210 0.146 0.234 0.258 0.227 0.382 0.154 0.145 0.325 0.099 0.390 0.342 0.422 0.387 0.477 0.763 0.457 0.739 0.742 0.753 0.525
Table B3: MAPE values of GLMs for each crop and region (Test Set)
Crop Region Model
@@
Corn (Grain) CEA 0.241 0.224 0.288 0.234 0.214 0.204 0.201 0.318 0.225 0.317 0.258 0.309 0.244 0.274 0.211 0.232 0.085 0.176 0.104 0.114 0.134 0.122
CWP 0.129 0.251 0.351 0.254 0.229 0.314 0.243 0.209 0.277 0.288 0.238 0.244 0.278 0.240 0.243 0.284 0.301 0.164 0.147 0.268 0.226 0.319
MID 0.281 0.329 0.301 0.321 0.251 0.314 0.255 0.340 0.350 0.305 0.367 0.292 0.304 0.293 0.261 0.275 0.293 0.279 0.261 0.275 0.347 0.279
SEA 0.168 0.263 0.399 0.269 0.243 0.330 0.238 0.310 0.302 0.339 0.337 0.262 0.227 0.235 0.222 0.246 0.323 0.234 0.347 0.312 0.146 0.321
SPL 0.159 0.232 0.250 0.244 0.187 0.264 0.184 0.151 0.241 0.248 0.190 0.168 0.235 0.208 0.233 0.181 0.261 0.229 0.168 0.246 0.068 0.291
SWP 0.176 0.110 0.214 0.124 0.108 0.120 0.051 0.157 0.107 0.225 0.136 0.124 0.102 0.057 0.069 0.102 0.058 0.072 0.049 0.063 0.159 0.077
Wheat (Winter) CEA 0.174 0.182 0.316 0.162 0.201 0.198 0.176 0.248 0.186 0.307 0.235 0.216 0.235 0.231 0.178 0.211 0.033 0.162 0.045 0.059 0.041 0.146
CWP 0.218 0.213 0.220 0.181 0.150 0.225 0.181 0.198 0.226 0.140 0.220 0.138 0.154 0.119 0.112 0.174 0.171 0.187 0.170 0.144 0.201 0.161
MID 0.254 0.290 0.283 0.266 0.242 0.269 0.248 0.317 0.310 0.285 0.325 0.293 0.302 0.291 0.282 0.273 0.291 0.274 0.273 0.282 0.274 0.307
SEA 0.159 0.220 0.346 0.158 0.233 0.269 0.233 0.247 0.225 0.295 0.269 0.274 0.286 0.276 0.165 0.267 0.106 0.216 0.072 0.081 0.085 0.107
SPL 0.177 0.212 0.210 0.183 0.155 0.205 0.175 0.146 0.213 0.214 0.172 0.184 0.213 0.195 0.208 0.197 0.127 0.210 0.111 0.110 0.154 0.093
SWP 0.144 0.114 0.061 0.133 0.073 0.067 0.063 0.167 0.086 0.071 0.124 0.121 0.037 0.073 0.051 0.055 0.107 0.043 0.055 0.068 0.078 0.220
Soybeans CEA 0.193 0.178 0.255 0.185 0.163 0.153 0.154 0.271 0.179 0.287 0.207 0.264 0.203 0.225 0.187 0.170 0.100 0.143 0.100 0.111 0.107 0.093
MID 0.232 0.283 0.280 0.273 0.215 0.275 0.223 0.307 0.304 0.279 0.318 0.268 0.266 0.270 0.240 0.239 0.240 0.243 0.270 0.239 0.243 0.308
SEA 0.231 0.289 0.319 0.294 0.231 0.276 0.238 0.319 0.309 0.285 0.324 0.263 0.218 0.236 0.221 0.244 0.093 0.229 0.060 0.088 0.106 0.216
SPL 0.184 0.226 0.229 0.221 0.157 0.250 0.175 0.175 0.232 0.231 0.183 0.166 0.212 0.184 0.195 0.153 0.108 0.196 0.114 0.103 0.146 0.129
[…]
Crop Region Model

### v2

@@ 5 Concluding Remarks
This study investigates the Actuaries Climate IndexTM within both the agricultural and financial sectors, comparing its effectiveness with widely recognized weather-based indexes. In the agricultural context, the analysis focuses on evaluating the predictive capabilities of ACI components and traditional WBIs for three major crops. By fitting 22 distinct models, each constructed with different combinations of climate indexes, the study examines the individual and joint contributions of each weather index to overall model performance. For the financial market application, the analysis centers on the payoffs of weather derivative contracts that use WBIs and ACI components as underlying variables. By assessing the strength of correlations among payoffs and comparing their ECDF values, the study evaluates the potential of these indexes for broader financial use.
Overall, based on the low MAPE values achieved in crop yield modeling and the consistency observed among the derivative payoffs, the results indicate that climate indexes demonstrate strong predictive performance and stable financial behavior. These findings highlight their applicability as effective climate risk indicators in both agricultural forecasting and weather-derivative markets. The key insights drawn from the analyses can be summarized as follows:
@@
    The four principal components derived through FPCA effectively capture the underlying patterns in crop yield variations.
@@
    The ACI components and their WBI counterparts exhibit strong correlation, and deliver similar predictive performance in predicting crop yields, especially with machine learning algorithms.
@@
    The inclusion of indexes representing additional weather conditions-such as wind speed, and sea-level changes-alongside temperature and precipitation, improves model performance.
@@
  6. (vi)
    The ACI components and their WBI counterparts generate similar derivative payoffs during comparable time periods, particularly when evaluated using the HBA method.
The outcomes suggest that the ACI framework holds strong potential as a comprehensive climate risk indicator, not only for the agricultural sector but also for the finance and insurance industries, where climate-related metrics are increasingly critical for risk assessment and decision-making.
Despite the promising performance metrics achieved, the study can be further improved in several aspects. One notable limitation is that regional crop yields are assessed by averaging the annual yields of the states within each region. This approach overlooks the fact that some states may have a greater influence on regional agricultural output due to their size or productivity levels. Moreover, the selection of the best-performing models for each crop is based on regional average results, which may not accurately reflect optimal performance at the regional or state level. Consequently, identifying and evaluating the best models for each specific region or state would enhance the precision and applicability of the findings, particularly for targeted applications in the insurance sector.
@@
From a financial perspective, the study further encourages the broader adoption of the Actuaries Climate IndexTM in risk management and hedging strategies. Demonstrating its applicability within weather-derivative framework underscores its potential utility for energy companies, insurers, and agribusinesses seeking to mitigate exposure to weather-related risks.
@@ Appendix A Preliminaries
In this section, an overview of Generalized Linear Models (GLMs) and Generalized Additive Models (GAMs), which belong to the class of generalized statistical models, as well as Extreme Gradient Boosting (XGB) and Light Gradient Boosting Machine (LGBM), which are classified as machine learning algorithms, is provided. Furthermore, the Principal Component Analysis (PCA) and Functional Principal Component Analysis (FPCA), two dimensionality reduction techniques used to preprocess and simplify data sets before being incorporated into the models, are reviewed.
@@ A.1 Generalized Statistical Models
where $\beta_{0}$ is the intercept term, and $f_{h}(X_{h})$ $(h=1,\dots,p)$ are smooth functions that model the non-linear effects of the predictors $X_{h}$ on the response variable. The flexibility of GAMs allow them to capture complex relationships while retaining interpretability, making it an effective tool for modeling real-world problems with non-linear dependencies.
@@ A.2 Machine Learning Algorithms
Here, $\ell$ denotes a differentiable loss function, $\hat{y}_{i}^{(t-1)}$ is the prediction at iteration t1t-1, $\Omega(f_{t})$ penalizes the complexity of each tree $f_{t}$, with $T$ being the number of leaves and $\mathbf{w}$ the vector of leaf weights.
@@ A.3 Dimensionality Reduction in Modeling
Given a data matrix $X$ of size $n\times p$ where $n$ represents the number of observations and $p$ is the number of variables, each element in $X$ can be standardized around its mean value, normalized by its standard deviation, thus obtaining a matrix of $Z$. PCA seeks to find the principal components by calculating the eigenvectors and eigenvalues of the covariance matrix, $\Sigma_{p\times p}$, given by:
@@
$\int_{a}^{b}C(s,t)\phi_{k}(t)dt=\lambda_{k}\phi_{k}(s).$ (24)
Appendix B Tables
@@
Group-1 CDD Mean ± SD 507 ± 98 245 ± 73 618 ± 109 1178 ± 110 795 ± 90 892 ± 105
Min. 326 85 356 929 527 697
Max. 704 428 839 1465 1017 1145
T90 Mean ± SD 0.44 ± 1.22 0.53 ± 1.33 0.01 ± 0.89 0.58 ± 1.23 0.58 ± 1.44 1.14 ± 1.78
Min. -1.48 -1.65 -1.31 -1.24 -1.61 -1.91
Max. 3.57 5.15 3.57 4.48 5.32 5.87
Group-2 HDD Mean ± SD 3266 ± 281 2723 ± 186 3647 ± 337 1874 ± 237 3339 ± 274 2372 ± 167
Min. 2715 2285 2901 1416 2797 2025
Max. 3832 3238 4382 2452 4209 2799
T10 Mean ± SD -0.49 ± 1.06 -0.17 ± 0.90 -0.32 ± 0.97 -0.47 ± 0.98 -0.30 ± 0.93 -0.33 ± 0.89
@@
Group-3 PRE Mean ± SD 11.35 ± 2.23 8.30 ± 1.86 10.12 ± 1.75 13.52 ± 2.61 7.04 ± 1.19 3.61 ± 1.04
Min. 6.80 3.88 6.11 7.50 4.49 1.55
Max. 19.04 13.50 13.82 20.38 10.21 7.02
P Mean ± SD 0.12 ± 1.01 0.60 ± 1.17 0.29 ± 1.05 -0.02 ± 0.93 0.35 ± 1.38 0.20 ± 0.97
Min. -2.06 -1.43 -2.13 -2.07 -1.65 -1.92
Max. 3.48 2.84 2.95 1.85 5.26 2.93
@@
Table B2: MAPE values of GLM for each crop, model, and region
Set Crop Region Model
@@
Train Corn CEA 0.345 0.343 0.350 0.344 0.347 0.344 0.344 0.346 0.343 0.347 0.344 0.349 0.345 0.343 0.344 0.343 0.341 0.344 0.342 0.342 0.343 0.342
CWP 0.388 0.391 0.389 0.390 0.388 0.391 0.390 0.391 0.390 0.397 0.391 0.396 0.397 0.397 0.397 0.397 0.396 0.398 0.396 0.397 0.393 0.393
MID 0.357 0.357 0.357 0.356 0.357 0.357 0.355 0.359 0.356 0.357 0.358 0.359 0.357 0.355 0.355 0.354 0.355 0.354 0.355 0.354 0.354 0.356
SEA 0.538 0.543 0.552 0.534 0.547 0.551 0.549 0.539 0.541 0.553 0.537 0.548 0.551 0.549 0.543 0.543 0.538 0.542 0.540 0.541 0.540 0.539
SPL 0.267 0.270 0.284 0.262 0.273 0.244 0.227 0.290 0.290 0.237 0.288 0.248 0.242 0.232 0.237 0.224 0.139 0.231 0.135 0.137 0.142 0.231
SWP 0.278 0.280 0.365 0.267 0.280 0.249 0.228 0.283 0.290 0.322 0.268 0.243 0.226 0.226 0.260 0.221 0.246 0.234 0.239 0.221 0.246 0.279
Wheat CEA 0.324 0.327 0.325 0.326 0.326 0.326 0.328 0.327 0.326 0.326 0.326 0.327 0.328 0.329 0.325 0.329 0.331 0.324 0.327 0.324 0.324 0.326
CWP 0.255 0.254 0.253 0.254 0.260 0.259 0.255 0.256 0.257 0.244 0.256 0.245 0.252 0.249 0.251 0.250 0.249 0.254 0.251 0.251 0.247 0.259
MID 0.336 0.334 0.334 0.335 0.335 0.334 0.334 0.334 0.335 0.333 0.335 0.335 0.335 0.334 0.333 0.334 0.334 0.334 0.333 0.334 0.334 0.335
SEA 0.342 0.341 0.343 0.341 0.342 0.347 0.344 0.343 0.341 0.342 0.343 0.343 0.342 0.343 0.343 0.344 0.340 0.344 0.340 0.342 0.342 0.343
SPL 0.214 0.222 0.218 0.216 0.218 0.220 0.220 0.222 0.219 0.219 0.222 0.221 0.218 0.218 0.215 0.219 0.215 0.219 0.215 0.217 0.215 0.224
SWP 0.224 0.219 0.226 0.220 0.209 0.207 0.206 0.224 0.219 0.213 0.218 0.195 0.200 0.203 0.208 0.206 0.211 0.207 0.208 0.209 0.211 0.223
Soybeans CEA 0.323 0.325 0.313 0.324 0.310 0.308 0.310 0.322 0.324 0.312 0.320 0.311 0.312 0.312 0.313 0.309 0.316 0.310 0.315 0.310 0.309 0.321
MID 0.293 0.293 0.293 0.292 0.294 0.294 0.294 0.294 0.293 0.294 0.294 0.294 0.294 0.293 0.293 0.292 0.293 0.292 0.293 0.292 0.292 0.291
SEA 0.326 0.323 0.330 0.325 0.329 0.329 0.328 0.327 0.324 0.329 0.325 0.328 0.328 0.328 0.326 0.326 0.323 0.326 0.324 0.325 0.322 0.321
SPL 0.248 0.240 0.243 0.240 0.245 0.243 0.241 0.243 0.245 0.240 0.240 0.241 0.241 0.241 0.242 0.240 0.240 0.243 0.243 0.240 0.240 0.239
Test Corn CEA 0.600 0.590 0.533 0.607 0.592 0.568 0.589 0.593 0.594 0.538 0.596 0.572 0.569 0.590 0.589 0.592 0.614 0.577 0.608 0.620 0.621 0.594
CWP 0.624 0.623 0.624 0.621 0.626 0.619 0.616 0.631 0.621 0.603 0.627 0.614 0.605 0.604 0.598 0.604 0.601 0.598 0.595 0.601 0.612 0.614
MID 0.575 0.588 0.578 0.594 0.590 0.585 0.599 0.577 0.592 0.575 0.595 0.570 0.586 0.581 0.581 0.584 0.581 0.585 0.581 0.584 0.585 0.590
SEA 0.800 0.780 0.746 0.810 0.780 0.761 0.782 0.810 0.792 0.744 0.808 0.779 0.762 0.774 0.787 0.787 0.805 0.785 0.798 0.792 0.793 0.804
SPL 0.228 0.269 0.287 0.263 0.236 0.273 0.235 0.246 0.268 0.254 0.257 0.243 0.258 0.240 0.248 0.223 0.172 0.246 0.166 0.179 0.177 0.147
SWP 0.170 0.169 0.300 0.163 0.147 0.178 0.158 0.177 0.183 0.294 0.182 0.167 0.171 0.148 0.153 0.207 0.155 0.201 0.164 0.225 0.246 0.172
Wheat CEA 0.540 0.535 0.553 0.536 0.552 0.541 0.542 0.533 0.536 0.550 0.531 0.538 0.541 0.541 0.565 0.542 0.538 0.546 0.551 0.546 0.529 0.532
CWP 0.439 0.425 0.418 0.411 0.416 0.403 0.408 0.439 0.418 0.379 0.420 0.387 0.371 0.361 0.375 0.380 0.374 0.379 0.374 0.373 0.396 0.407
MID 0.532 0.541 0.562 0.529 0.553 0.556 0.554 0.534 0.546 0.561 0.534 0.545 0.558 0.552 0.561 0.554 0.552 0.556 0.561 0.554 0.556 0.539
SEA 0.605 0.603 0.572 0.614 0.615 0.596 0.610 0.602 0.595 0.581 0.598 0.614 0.596 0.603 0.603 0.600 0.630 0.603 0.626 0.610 0.616 0.601
SPL 0.366 0.377 0.363 0.375 0.360 0.363 0.370 0.378 0.377 0.365 0.379 0.371 0.372 0.377 0.389 0.385 0.370 0.384 0.365 0.366 0.363 0.363
SWP 0.220 0.241 0.276 0.232 0.213 0.235 0.218 0.228 0.254 0.268 0.233 0.214 0.247 0.231 0.231 0.252 0.216 0.247 0.223 0.253 0.246 0.225
Soybeans CEA 0.513 0.506 0.477 0.516 0.500 0.486 0.501 0.516 0.514 0.481 0.514 0.495 0.492 0.508 0.547 0.508 0.526 0.530 0.550 0.544 0.560 0.507
MID 0.481 0.494 0.496 0.493 0.489 0.490 0.493 0.495 0.496 0.489 0.498 0.485 0.489 0.486 0.488 0.485 0.486 0.488 0.488 0.485 0.488 0.505
SEA 0.584 0.573 0.504 0.604 0.548 0.527 0.555 0.601 0.567 0.509 0.588 0.556 0.527 0.548 0.553 0.555 0.621 0.545 0.610 0.625 0.627 0.615
SPL 0.443 0.436 0.430 0.436 0.426 0.434 0.412 0.454 0.436 0.419 0.447 0.422 0.414 0.405 0.419 0.404 0.444 0.420 0.446 0.446 0.433 0.445
Table B3: MAPE values of GAM for each crop, model, and region
Set Crop Region Model
@@
Train Corn CEA 0.345 0.342 0.350 0.343 0.347 0.347 0.344 0.344 0.343 0.347 0.343 0.348 0.345 0.345 0.343 0.343 0.342 0.345 0.342 0.343 0.343 0.342
CWP 0.388 0.393 0.396 0.391 0.393 0.395 0.392 0.391 0.394 0.399 0.392 0.397 0.399 0.397 0.397 0.398 0.397 0.399 0.398 0.398 0.396 0.396
MID 0.357 0.355 0.360 0.355 0.357 0.359 0.356 0.358 0.355 0.357 0.357 0.358 0.358 0.355 0.355 0.354 0.355 0.354 0.355 0.354 0.354 0.355
SEA 0.539 0.544 0.552 0.538 0.544 0.548 0.548 0.541 0.541 0.550 0.538 0.547 0.550 0.547 0.540 0.543 0.537 0.541 0.539 0.539 0.542 0.540
SPL 0.204 0.261 0.252 0.256 0.218 0.220 0.225 0.257 0.219 0.187 0.233 0.233 0.220 0.196 0.208 0.220 0.109 0.228 0.073 0.107 0.119 0.188
SWP 0.236 0.250 0.336 0.260 0.212 0.174 0.225 0.215 0.277 0.270 0.151 0.186 0.212 0.192 0.255 0.189 0.213 0.174 0.167 0.143 0.130 0.229
Wheat CEA 0.327 0.327 0.326 0.327 0.327 0.326 0.328 0.328 0.327 0.327 0.327 0.329 0.328 0.330 0.325 0.330 0.328 0.327 0.328 0.326 0.325 0.326
CWP 0.242 0.238 0.254 0.256 0.251 0.258 0.249 0.250 0.249 0.246 0.254 0.244 0.250 0.241 0.250 0.251 0.248 0.251 0.251 0.251 0.248 0.250
MID 0.335 0.334 0.334 0.334 0.334 0.335 0.333 0.334 0.334 0.333 0.335 0.335 0.334 0.333 0.332 0.335 0.333 0.334 0.332 0.335 0.334 0.335
SEA 0.344 0.343 0.343 0.341 0.342 0.346 0.346 0.345 0.341 0.342 0.345 0.344 0.343 0.346 0.347 0.345 0.348 0.346 0.345 0.344 0.342 0.345
SPL 0.214 0.222 0.218 0.217 0.217 0.221 0.222 0.221 0.218 0.220 0.224 0.222 0.218 0.220 0.214 0.219 0.216 0.219 0.217 0.216 0.215 0.224
SWP 0.209 0.193 0.210 0.200 0.198 0.198 0.184 0.219 0.216 0.210 0.205 0.175 0.189 0.182 0.191 0.174 0.209 0.184 0.207 0.198 0.190 0.223
Soybeans CEA 0.324 0.325 0.312 0.322 0.308 0.304 0.305 0.322 0.321 0.307 0.318 0.309 0.311 0.309 0.308 0.306 0.314 0.307 0.314 0.307 0.305 0.313
MID 0.293 0.293 0.293 0.292 0.294 0.294 0.293 0.294 0.292 0.294 0.293 0.293 0.294 0.293 0.293 0.292 0.293 0.291 0.293 0.292 0.291 0.292
SEA 0.327 0.323 0.330 0.325 0.328 0.329 0.327 0.327 0.325 0.328 0.325 0.329 0.328 0.327 0.326 0.325 0.323 0.325 0.324 0.325 0.323 0.321
SPL 0.248 0.245 0.248 0.245 0.250 0.247 0.246 0.244 0.245 0.240 0.242 0.246 0.246 0.248 0.244 0.243 0.242 0.243 0.243 0.242 0.242 0.242
Test Corn CEA 0.608 0.591 0.532 0.615 0.592 0.571 0.594 0.599 0.596 0.538 0.595 0.598 0.562 0.568 0.602 0.646 0.621 0.576 0.609 0.618 0.615 0.715
CWP 0.624 0.622 0.611 0.614 0.623 0.618 0.612 0.632 0.617 0.598 0.630 0.614 0.607 0.611 0.593 0.600 0.596 0.597 0.606 0.603 0.594 0.619
MID 0.585 0.588 0.565 0.591 0.598 0.587 0.597 0.602 0.594 0.583 0.602 0.588 0.581 0.587 0.618 0.587 0.587 0.581 0.618 0.587 0.581 0.593
SEA 0.805 0.779 0.746 0.811 0.779 0.750 0.776 0.813 0.791 0.739 0.804 0.773 0.763 0.779 0.790 0.783 0.796 0.768 0.753 0.767 0.782 0.796
SPL 0.233 0.261 0.307 0.257 0.312 0.275 0.230 0.254 0.267 0.295 0.266 0.246 0.295 0.283 0.257 0.224 0.103 0.242 0.225 0.110 0.121 0.119
SWP 0.210 0.169 0.359 0.162 0.200 0.208 0.162 0.228 0.185 0.281 0.290 0.249 0.185 0.173 0.159 0.211 0.124 0.210 0.119 0.168 0.235 0.264
Wheat CEA 0.605 0.535 0.542 0.519 0.550 0.543 0.546 0.533 0.534 0.546 0.530 0.539 0.540 0.548 0.590 0.552 0.535 0.547 0.558 0.542 0.525 0.538
CWP 0.423 0.408 0.416 0.393 0.423 0.401 0.397 0.435 0.409 0.388 0.416 0.395 0.365 0.366 0.373 0.384 0.374 0.391 0.370 0.364 0.396 0.401
MID 0.524 0.539 0.564 0.529 0.542 0.562 0.555 0.543 0.545 0.561 0.528 0.542 0.560 0.557 0.568 0.562 0.557 0.560 0.568 0.562 0.560 0.539
SEA 0.628 0.606 0.567 0.616 0.612 0.590 0.608 0.609 0.606 0.580 0.607 0.596 0.586 0.617 0.595 0.596 0.576 0.601 0.613 0.871 0.614 0.596
SPL 0.371 0.381 0.363 0.386 0.342 0.371 0.370 0.385 0.381 0.376 0.370 0.363 0.376 0.384 0.392 0.381 0.364 0.389 0.367 0.366 0.408 0.365
SWP 0.230 0.237 0.273 0.235 0.202 0.230 0.247 0.225 0.255 0.265 0.227 0.216 0.244 0.216 0.231 0.278 0.226 0.271 0.233 0.255 0.246 0.224
Soybeans CEA 0.517 0.505 0.478 0.521 0.502 0.500 0.521 0.516 0.509 0.487 0.504 0.508 0.528 0.497 0.561 0.537 0.520 0.556 0.552 0.540 0.547 0.520
MID 0.482 0.494 0.498 0.492 0.473 0.495 0.495 0.492 0.495 0.486 0.498 0.485 0.488 0.486 0.487 0.480 0.486 0.485 0.487 0.480 0.485 0.507
SEA 0.580 0.571 0.491 0.609 0.539 0.518 0.540 0.611 0.580 0.495 0.590 0.561 0.532 0.565 0.528 0.564 0.593 0.577 0.577 0.562 0.599 0.600
SPL 0.455 0.451 0.440 0.444 0.439 0.434 0.418 0.462 0.431 0.429 0.469 0.471 0.418 0.432 0.415 0.399 0.442 0.417 0.441 0.453 0.433 0.461
Table B4: MAPE values of XGB for each crop, model, and region
Set Crop Region Model